an 79.0 kg athlete running a "4-minute mile" (i.e. 4.00 min/mile) . Answer in nm.
1
Expert's answer
2019-09-30T04:54:52-0400

λ = h / p = h / m x v=6.62 x 10 -34 /79000 x 107.2896=

7.81039992e-41 m.

7.81039992e-32 nm